# Integration vs segregation: evidence from measuring and comparing interdisciplinarity in funded and unfunded research using a network approach

### v2

@@ 3.2 Network Measures for ID
where $m$ is the sum of link weights, $s_{i}$ refers to the strength of node $i$, and $\frac{s_{i}s_{j}}{2m}$ is the expected link strength between $i$ and $j$ assuming a random distribution of connections which preserves the strength distribution across nodes. A high modularity $Q$ represents a well-defined community structure with many intra-community links and few links connecting separate communities while a low $Q$ indicates a weak community structure.
@@ 4 Results
4.1 Identification of temporal regimes in interdisciplinary research
@@
4.2 Trends in compartmentalisation and integration of funded and unfunded research
@@
From the above, we infer that knowledge integration in funded research tends to reinforce community structure. This indicates that funded research tends to be more specialised and stays on the conservative side, i.e., it focuses on deepening already established relationships of diseases. Knowledge integration in unfunded research, on the other hand, tends to take place through weakening community structures, which indicates that unfunded research is relatively less conservative and focuses more on bridging distant diseases. These interpretations are further supported by results shown in Figure 3 and Figure 4, where we have compared relative increments in link strengths between time periods for the strongest (top 5%) and weaker (bottom 50%) links in the evolution of IDR in funded and unfunded research. We note, that the strongest links in funded research have consistently gained more in link strength compared to unfunded research (relative gains $4.96>3.74$ from 1995-2008 to 2009-2015 and even more so $4.13>2.67$ from 2009-2015 to 2016-2022). In contrast, for the majority of weaker links, links have gained more in strength for unfunded research than for funded research (relative gains $0.26>0.16$ from 1995-2008 to 2009-2015 and $0.22>0.15$ from 2009-2015 to 2016-2022). Both observations make it very clear that strength gains in funded research tend to be more aligned with already strong connections, whereas strength gains in unfunded research tend to be more exploratory, reinforcing weak connections.
@@
4.3 Analysing research funding into Interdisciplinary Research
@@
4.4 Comparing the roles of diseases in knowledge integration
To identify the diseases with the strongest global and local impact, we ranked the diseases by strength and betweenness. Table 2 reports the top three ranked strengths $s_{i}$ and betweenness centrality $b_{i}$ of the infectious diseases for F and U regimes. $s_{i}$ represents the knowledge integration at a local level and $b_{i}$ at a global level, thus comparing these two allows us to compare the local and global impact of the top-ranked diseases. Comparing the different time regimes we make the following observations. Regarding the betweennees centrality $b_{i}$, we find HIV has the highest $b_{i}$ in all regimes for both F and U, followed by other prominent diseases like tuberculosis, HBV, or malaria. This indicates that prominent infectious diseases like HIV stay in the centre of the network and bridge distant bodies of knowledge. Regarding $s_{i}$, we find DTP are the top three in $s_{i}$ in F2, F3, U2 and U3 with a remarkable increase in magnitude through time, but none of DTP appears in the top three $b_{i}$ of F or U regimes. This indicates that DTP stay on the periphery of the network and only enhance established connections between bodies of knowledge.
@@ 4.4 Comparing the roles of diseases in knowledge integration
Table 2: Top three ranked diseases, ranked by node strength $s_{i}$ and betweenness centrality $b_{i}$ for the funded (F) and unfunded (U) regimes.
@@
4.5 Quantifying the Role of Coronavirus Research in IDR
@@ 5 Discussion
We summarise the results and discuss the practical implications of this study in Section 5.1, then discuss real-world factors behind the results, limitations and future work in Section 5.2.
5.1 Summary of results and practical implications
@@
The results of this study contribute to the understanding of the roles played by the research on individual infectious disease in interdisciplinary knowledge generation, and the relationship between global public health emergencies and interdisciplinary research efforts. These could provide valuable insights for future health priority setting and resource allocation.
The proposed analytical framework of funded and unfunded IDR can be applied to any field of science. It could provide valuable insights to researchers, funders and policymakers on the evolution of IDR and the over/under-funded interdisciplinary areas, supporting funding decisions to facilitate IDR.
5.2 Real-world factors, limitations, and future work
The observed compartmentalisation in funded research output in terms of ID is consistent with the lack of appreciation towards IDR during research evaluations (Bromham et al., 2016; Boudreau et al., 2016) and the observed conservativeness of research funding (Packalen and Bhattacharya, 2020). Funding sources, especially those that routinely issue topic-focused calls, seem to have the greatest influence on most science and health research which are resource intensive (Thelwall et al., 2023; Whitley et al., 2018); however, the degree of such influence depends on individual funders’ requirements (Serrano Velarde, 2018; Luukkonen and Thomas, 2016; Thelwall et al., 2023). The observation that unfunded research is more widely exploratory in terms of ID is consistent with the fact that unfunded researchers have a higher degree of freedom in setting their own goals (Edwards, 2022).
@@ 5.2 Real-world factors, limitations, and future work
The rising consistency between funded and unfunded research might suggest a better allocation of funding in the sense that interdisciplinary research areas with high scientific interest have been addressed by research funding. However, this would need further in-depth examination as we did not take into account the complex interplay between funded and unfunded research in our study. We want to also note that the focus of our work is on topic-level ID instead of ID based on conventional classification systems. The granularity of the choice of classifications is likely to have an impact on the resulting ID (Rousseau et al., 2019). We leave in-depth examinations of such impact for future work.
@@
We distinguished (the research on) infectious diseases with two different patterns of knowledge integration: diseases that tend to integrate locally (e.g. DTP) and diseases that tend to integrate globally (e.g. HIV, tuberculosis, and malaria). There is no simple answer to what drives these integrations as the collection of knowledge on infectious diseases spans multiple domains including pathogenesis, diagnosis, cause, treatment, prognosis, spread, prevention, social impact and policy. However, there may be fundamentally two types of driving factors: intentional or unintentional (Wagner et al., 2011; Glänzel and Debackere, 2022). Policymakers, funders and researchers have been targeting specific groups of diseases to meet their respective priorities and goals: these are intentional efforts, and they tend to focus on more established groups of diseases that have been associated with great disease burdens at that time. For instance, the WHO has created global health programmes focusing on Hepatitis B and C [^16] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/hepatitis/overview (Date last accessed: 2024.9.4), STIs [^17] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/stis/overview (Date last accessed: 2024.9.4), and vector-borne infections [^18] https://www.who.int/publications/i/item/9789241512978 (Date last accessed: 2024.9.4). Therefore, it might be that emergence of the diseases with strong local integration is the result of such intentional effort. For example, DTP, the diseases found with the largest extent of local integration, have been a part of the WHO Expanded Programme on Immunization [^19] https://www.who.int/teams/immunization-vaccines-and-biologicals/essential-programme-on-immunization (Date last accessed: 2024.9.4), and a focus of Gavi, the Vaccine Alliance [^20] https://www.gavi.org/news/media-room/gavi-helps-dtp3-coverage-rise-after-stagnation (Date last accessed: 2024.9.4), and there has been ongoing development of a variety of DTP-related combination vaccines [^21] https://www.who.int/teams/health-product-policy-and-standards/standards-and-specifications/norms-and-standards/vaccine-standardization/dt-based-combined-vaccines (Date last accessed: 2024.9.4).

### v3

@@ 3.2 Network measures and analysis
Closeness centrality (Freeman et al., 2002) offers a comparative perspective to betweenness centrality. As a node-level measure, it quantifies how close a disease is to all other diseases in the network. It is computed as
$c_{i}=\frac{1}{\frac{1}{n}\sum_{j=1}^{n}d(i,j)}$ (4)
where $n$ is the number of all reachable node from $i$.
@@
where $m$ is the sum of link weights, $s_{i}$ refers to the strength of node $i$, and $\frac{s_{i}s_{j}}{2m}$ is the expected link strength between $i$ and $j$ assuming a random distribution of connections which preserves the strength distribution across nodes. A high modularity $Q$ represents a well-defined community structure with many intra-community links and few links connecting separate communities, i.e., compartmentalisation (Cohen and D’Esposito, 2016), while a low $Q$ indicates a weak community structure.
@@ 4 Results
4.1 RQ1: Identification of temporal regimes in interdisciplinary research
@@
4.2 RQ2: Compartmentalisation and integration of funded and unfunded research
@@
From the above, we infer that knowledge integration in funded research tends to reinforce community structure. This indicates that funded research tends to be more specialised and stays on the conservative side, i.e., it focuses on deepening already established relationships of diseases (Rzhetsky et al., 2015; Foster et al., 2015). Knowledge integration in unfunded research, on the other hand, tends to take place through weakening community structures, which indicates that unfunded research is relatively less conservative and focuses more on bridging distant diseases [^16] The observation that unfunded research is more widely exploratory in terms of ID is consistent with the fact that unfunded researchers have a higher degree of freedom in setting their own goals (Edwards, 2022).. These interpretations are further supported by results shown in Figure 5 and Figure 6, where we have compared relative increments in link strengths between time periods for the strongest (top 5%) and weaker (bottom 50%) links in the evolution of IDR in funded and unfunded research. We note, that the strongest links in funded research have consistently gained more in link strength compared to unfunded research (relative gains $4.96>3.74$ from 1995-2008 to 2009-2015 and even more so $4.13>2.67$ from 2009-2015 to 2016-2022). In contrast, for the majority of weaker links, links have gained more in strength for unfunded research than for funded research (relative gains $0.26>0.16$ from 1995-2008 to 2009-2015 and $0.22>0.15$ from 2009-2015 to 2016-2022). Both observations make it very clear that strength gains in funded research tend to be more aligned with already strong connections, whereas strength gains in unfunded research tend to be more exploratory, reinforcing weak connections.
@@
4.3 RQ3: Analysing research funding into Interdisciplinary Research
@@
4.4 RQ4: Comparing the roles of diseases in knowledge integration
To identify the diseases with the strongest global and local impact, we ranked the diseases by strength, betweenness and closeness. Table 2 reports the top three ranked strengths $s_{i}$, betweenness centrality $b_{i}$ and closeness centrality $c_{i}$ of the infectious diseases for F and U regimes. $s_{i}$ represents the knowledge integration at a local level and $b_{i}$ and $c_{i}$ at a global level, thus comparing these three allows us to compare the local and global impact of the top-ranked diseases. Comparing the different time regimes we make the following observations. Regarding the betweennees centrality $b_{i}$ and closeness centrality $c_{i}$, we find HIV has the highest $b_{i}$ and $c_{i}$ in all regimes for both F and U, followed by other prominent diseases like tuberculosis, HBV, or malaria. This indicates that prominent infectious diseases like HIV stay in the centre of the network and bridge distant bodies of knowledge. Regarding $s_{i}$, we find DTP are the top three in $s_{i}$ in F2, F3, U2 and U3 with a remarkable increase in magnitude through time, but none of DTP appears in the top three $b_{i}$ of F or U regimes. This indicates that DTP stay on the periphery of the network and only enhance established connections between bodies of knowledge.
@@ 4.4 RQ4: Comparing the roles of diseases in knowledge integration
Top Three $c_{i}$Centerstack HIV(0.017)
HBV(0.016)
[…]
TB (0.019)
[…]
Measles(0.022)
Diph(0.022)
Table 2: Top three ranked diseases, ranked by node strength $s_{i}$, betweenness centrality $b_{i}$, and closeness centrality $c_{i}$ for the funded (F) and unfunded (U) regimes.
@@
4.5 RQ5: Quantifying the Role of Coronavirus Research in IDR
@@ 5 Discussion
We summarise the results and discuss the contribution of this study in Section 5.1, then discuss the implications of the results, limitations and future work in Section 5.2.
5.1 Summary of results and contributions
@@
The results of this study contribute to the understanding of the roles played by the research on individual infectious disease in interdisciplinary knowledge generation, and the relationship between global public health emergencies and interdisciplinary research efforts. These could provide valuable insights for future health priority setting, for example with horizon scanning for new and emerging threats to health, such as pandemic planning.
Overall, this research presents a generalisable framework to examine the impact of funding in interdisciplinary knowledge creation and provides important insights to science policy.
5.2 Implications, limitations and future work
Conservatism in scientific knowledge (Foster et al., 2015; Rzhetsky et al., 2015) manifests in infectious disease research through a compartmentalized structure shaped by funding patterns. This trend aligns with broader observations by Park et al. (2023) that scientific disruptiveness and innovation are declining due to an over-reliance on a narrower scope of existing knowledge—a ”conservative trap” (Fortunato et al., 2018). The conservatism in idea selection in science comes down to scientists’ inclination to prefer productivity over riskier innovation (IDR for instance), where the rewards of additional impact do not compensate for the risk of publishing nothing (Fortunato et al., 2018). Research institutions and funding agencies should encourage intelligent risk-taking through establishing mechanisms that diversify risks across a portfolio of scientific projects during evaluation (Rzhetsky et al., 2015), perhaps adopting models like the group-based evaluations at Bell Labs or the people-centred approach of the Howard Hughes Medical Institute.
Such conservatism could also be attributed to and reinforced by rigid disciplinary-based research evaluations (Rousseau et al., 2019; Woelert and Millar, 2013; Rylance, 2015; Fontana et al., 2022) where evaluators that are subject to bounded rationality and hold disciplinary-based standards tend to penalise novelty (Boudreau et al., 2016; Packalen and Bhattacharya, 2020; Woelert and Millar, 2013). As a consequence, despite being repeatedly advocated in science policy documents (Woelert and Millar, 2013), IDR becomes systematically disadvantaged (Bromham et al., 2016), forming the ”paradox of interdisciplinarity” (Woelert and Millar, 2013). Addressing this might require a dual approach: relaxing stringent classification systems, despite the political and bureaucratic costs (Woelert and Millar, 2013), and directly channelling support to interdisciplinary initiatives (Lyall et al., 2013) through schemes like the UKRI’s cross research council responsive mode pilot scheme or the Gates Foundation’s Grand Challenges. Moreover, establishing interdisciplinary research centres, as the NSF has done, could further support this shift, although the impact of such centres should be continuously assessed to confirm their efficacy (Woelert and Millar, 2013).
@@ 5.2 Implications, limitations and future work
The rising consistency between funded and unfunded research might suggest a better allocation of funding in the sense that interdisciplinary research areas with high scientific interest have been addressed by research funding. However, this would need further in-depth examination as we did not take into account the complex interplay between funded and unfunded research in our study.
We note that the focus of our work is on topic-level ID instead of ID based on conventional classification systems. Looking at a specific field like infectious disease research and considering the sub-topics allows us to capture the more granular level dynamics of knowledge integration that has typically been overlooked by past research (Rafols and Meyer, 2010; Rafols, 2014). However, we note that the granularity of the choice of classifications is likely to have an impact on the resulting ID (Rousseau et al., 2019) and we leave an in-depth examinations of the effects of granularity for future work. We also note that although this research focuses on the dynamics of infectious disease research, the proposed framework is generalisable to any other research field.
@@
We distinguished (the research on) infectious diseases with two different patterns of knowledge integration: diseases that tend to integrate locally (e.g. DTP) and diseases that tend to integrate globally (e.g. HIV, tuberculosis, and malaria). There is no simple answer to what drives these integrations as the collection of knowledge on infectious diseases spans multiple domains including pathogenesis, diagnosis, cause, treatment, prognosis, spread, prevention, social impact and policy (Hamburg, 2008). However, there may be fundamentally two types of driving factors: intentional or unintentional (Wagner et al., 2011; Glänzel and Debackere, 2022). Policymakers, funders and researchers have been targeting specific groups of diseases to meet their respective priorities and goals: these are intentional efforts, and they may tend to focus on more established groups of diseases that have been associated with great disease burdens at that time. For instance, the WHO has created global health programmes focusing on Hepatitis B and C [^17] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/hepatitis/overview (Date last accessed: 2024.9.4), STIs [^18] https://www.who.int/teams/global-hiv-hepatitis-and-stis-programmes/stis/overview (Date last accessed: 2024.9.4), and vector-borne infections [^19] https://www.who.int/publications/i/item/9789241512978 (Date last accessed: 2024.9.4). Therefore, it might be that emergence of the diseases with strong local integration is the result of such intentional effort. For example, DTP, the diseases found with the largest extent of local integration, have been a part of the WHO Expanded Programme on Immunization [^20] https://www.who.int/teams/immunization-vaccines-and-biologicals/essential-programme-on-immunization (Date last accessed: 2024.9.4), and a focus of Gavi, the Vaccine Alliance [^21] https://www.gavi.org/news/media-room/gavi-helps-dtp3-coverage-rise-after-stagnation (Date last accessed: 2024.9.4), and there has been ongoing development of a variety of DTP-related combination vaccines [^22] https://www.who.int/teams/health-product-policy-and-standards/standards-and-specifications/norms-and-standards/vaccine-standardization/dt-based-combined-vaccines (Date last accessed: 2024.9.4).
@@
6 Conclusion
This research provides a generalisable framework to examine the impact of funding in interdisciplinary knowledge creation. There is limited money available for global health research and development, especially from public and charitable funders. Thus, we must invest wisely. Effective research funding is vital in driving breakthroughs in science and technology. However, we caution that funded research output has witnessed a growing conservatism in the past decades, potentially slowing down scientific progress. We urge funding agencies and policymakers to better recognize and reward interdisciplinary contributions. Funding agencies and research institutions should prioritize mechanisms that encourage intelligent risk-taking.
@@ 6 Conclusion
Pair False Positive Rate
HIV-TB 2%
Dengue-Zika 1%
Tetanus-Diphtheria 0%
HCV-HBV 2%
Clamydia-Gonorreahea 3%
Table 1: False positive rate of knowledge integration in a paper’s abstract considering five representative pairs of infectious diseases. The sample size for each pair is 100. The false positive rates are consistently less than 3% and we found no bias towards particular disease pairs.